A flywheel has an angular momentum of 40 kg m^2 s^-1. The radius of gyration of the flywheel is 0.65m and its rotating at 250 RPM. 

Whats the angular velocity in rad/s?

Whats its mass?


1
Expert's answer
2021-06-21T11:41:54-0400
"\omega=250\frac{2\pi}{60}=26.18\frac{rad}{s}
\\m=\frac{L}{R_g^2\omega}
\\m=\frac{40}{0.65^2(26.18)}=3.6\ kg"